OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (zzr0ck3r):

what is the derivative?

OpenStudy (zzr0ck3r):

@emersonba123 ?

OpenStudy (anonymous):

That's all that the problem gave me. Do you need the equation? f(x+h)-f(x)/h

OpenStudy (zzr0ck3r):

yes

OpenStudy (zzr0ck3r):

do that and tell me what you get

OpenStudy (anonymous):

Do I substitute s(t) for f(x)?

OpenStudy (zzr0ck3r):

yep

OpenStudy (zzr0ck3r):

(-8 - 9(t+h)-(-8 - 9t))/h

OpenStudy (zzr0ck3r):

simplify and run the limit and tell me what you have

OpenStudy (anonymous):

I got -8

OpenStudy (zzr0ck3r):

should be -9

OpenStudy (precal):

isn't this just taking the derivative of s(t)

OpenStudy (zzr0ck3r):

yes they don't know the shortcut rules yet

OpenStudy (zzr0ck3r):

so from the deffinition

OpenStudy (precal):

really? I am sure they have seen the power rule at this point

OpenStudy (zzr0ck3r):

are you in the class?

OpenStudy (zzr0ck3r):

This is exactly the sort of problem they give you with the definition.

OpenStudy (zzr0ck3r):

(-8 - 9(t+h)-(-8 - 9t))/h = (-8-9t-9h+8+9t)/h = -9h/h = -9

OpenStudy (anonymous):

Okay guys. Pause. When I distribute, I get (-8-9t-9h+8+9t)/h and the -9t and +9t would cancel out and leave you with -8-9h+8/h

OpenStudy (zzr0ck3r):

so your derivative function is f(x) = -9

OpenStudy (anonymous):

and the 8's cancel out. Okay, I see it now.

OpenStudy (zzr0ck3r):

so what is f(1) = -9

OpenStudy (anonymous):

Wouldn't it just be -9 since there are no x's

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

yay! Thank you so much!

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

in about 1 week you will be able to answer this problem without doing anything, just by looking at it

OpenStudy (zzr0ck3r):

they will teach you a "trick" no more f(x+h).....

OpenStudy (anonymous):

d/dx cx^n = ncx^(n-1) Best rule ever!

OpenStudy (zzr0ck3r):

don't use it though, until your teacher tells you the rule.

OpenStudy (zzr0ck3r):

great for checking work though:)

OpenStudy (anonymous):

No just use it right away. Your teacher will be impressed and let you test out of the class

OpenStudy (zzr0ck3r):

lol

OpenStudy (zzr0ck3r):

also tell your teacher what he calls a linear function is actually an affine function, they will love you

OpenStudy (anonymous):

I have no idea what any of that is..

OpenStudy (zzr0ck3r):

ignore us @emersonba123

OpenStudy (anonymous):